OpenStudy (anonymous):

What is the slope intercept form that passes through the given point and is perpendicular (3,2), y=x+5? What is the answer

OpenStudy (anonymous):

In y=x+5, what is the slope?

OpenStudy (anonymous):

Y=-1x+1

OpenStudy (anonymous):

Best answer gets medal

OpenStudy (anonymous):

The slope is 1. So, Perp. slope would be (-1/1) or -1. y-y1=m(x-x1), with point (3,2)

OpenStudy (anonymous):

y-2=-1(x-3) solve for y

OpenStudy (anonymous):

Y=1

OpenStudy (anonymous):

y-2=-1(x-3) y-2=-x+3 +2 +2 y=-x+5

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

or you can write it as y=5-x

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

lets say we have y-5=-4(x+2) solve for y, don't forget the x

OpenStudy (anonymous):

leave the left side alone, so it stays as y-5 = .... distribute the -4 to x and 2